# Neutron skins probed in proton knockout from neutron-rich nuclei

### v1

@@ III.1 Numerical Results
where $A_{f}=A-1$ for (p,2p) and $A_{f}=A-2$ for (p,3p) reactions. Our model provides a microscopically motivated extension of the physical ideas underlying the Goldhaber model Goldhaber1974 ; PhysRevC.39.460 . In the latter, the momentum width arises from the vector sum of randomly oriented Fermi momenta, yielding $\Delta_{\parallel}^{2}\propto\rho\,(3p_{F}^{2}/5)$, with a constant bulk Fermi momentum $p_{F}$. In contrast, the present calculation naturally incorporates the surface bias of proton removal reactions and the reduction of the local Fermi momentum in low-density regions. Although only protons are removed, the neutron distribution modifies the geometry of the interaction region and shifts the sampling of proton momenta toward regions of lower density. This results in a systematic reduction of $\Delta_{\parallel}$ with increasing neutron skin thickness within the present reaction framework. The effect is more pronounced for the $(p,3p)$ channel, which samples the surface more strongly and involves a larger recoil.
@@ IV Conclusions
The results presented in this work indicate that inclusive (p,2p) and (p,3p) reactions exhibit a systematic sensitivity to neutron-skin thickness through both total cross sections and longitudinal momentum distributions. The decrease of cross sections with increasing neutron excess observed along the tin isotopic chain arises primarily from enhanced attenuation of the incoming and outgoing protons by the neutron-rich surface region. Because neutron-proton cross sections exceed proton-proton cross sections at intermediate and relativistic energies, the neutron density plays a dominant role in shaping the reaction transparency. As a result, nuclei with thicker neutron skins exhibit reduced effective access to interior protons, leading to smaller knockout cross sections.
@@ Acknowledgments
This work was supported by U.S. Department of Energy Office of Nuclear Physics under Contract No. DE-SC0026074 with East Texas A&M University.

### v2

@@ III.1 Numerical Results
where $A_{f}=A-1$ for (p,2p) and $A_{f}=A-2$ for (p,3p) reactions. Our model provides a microscopically motivated extension of the physical ideas underlying the Goldhaber model Goldhaber1974 ; PhysRevC.39.460 . In Eq. (23), $p_{0}$ is the Goldhaber width parameter characterizing the effective Fermi momentum of nucleons in the statistical fragmentation model. Throughout this work we adopt the standard value $p_{0}=90$ MeV/$c$, commonly used in analyses of relativistic fragmentation data. In the latter, the momentum width arises from the vector sum of randomly oriented Fermi momenta, yielding $\Delta_{\parallel}^{2}\propto\rho\,(3p_{F}^{2}/5)$, with a constant bulk Fermi momentum $p_{F}$. In contrast, the present calculation naturally incorporates the surface bias of proton removal reactions and the reduction of the local Fermi momentum in low-density regions. Although only protons are removed, the neutron distribution modifies the geometry of the interaction region and shifts the sampling of proton momenta toward regions of lower density. This results in a systematic reduction of $\Delta_{\parallel}$ with increasing neutron skin thickness within the present reaction framework. The effect is more pronounced for the $(p,3p)$ channel, which samples the surface more strongly and involves a larger recoil.
@@
Benchmark comparisons of calculated longitudinal momentum widths with experimental data for light nuclei such as 12C and 16O, where the number of active orbitals is small, would provide a valuable test of the present framework. Such an analysis is beyond the scope of the present work and will be addressed in future studies.
@@ IV Conclusions
The results presented in this work indicate that inclusive (p,2p) and (p,3p) reactions exhibit a systematic sensitivity to neutron-skin thickness through both total cross sections and longitudinal momentum distributions. The decrease of cross sections with increasing neutron excess observed along the tin isotopic chain arises primarily from enhanced attenuation of the incoming and outgoing protons by the neutron-rich surface region. Because neutron-proton cross sections exceed proton-proton cross sections at intermediate and relativistic energies, the neutron density plays a dominant role in shaping the reaction transparency. As a result, nuclei with thicker neutron skins exhibit reduced effective access to interior protons, leading to smaller knockout cross sections. We have clarified that while (p,2p) reactions do probe the nuclear interior, strong attenuation leads to effective surface weighting, particularly in neutron-rich nuclei. This clarification reconciles the present discussion with Refs. aumann:2013:PRC ; Bertulani.104.L061602 .
@@ Acknowledgments
C.A.B. is supported by U.S. Department of Energy Office of Nuclear Physics under Contract No. DE-SC0026074 with East Texas A&M University. R.V.L. is supported by INCT-FNA (Instituto Nacional de Ciência e Tecnologia, Física Nuclear e Aplicações), research Project No. 464898/2014-5, and also thanks to CAPES/CNPq.